OpenStudy (anonymous):

find the square root of 0? show work

OpenStudy (anonymous):

lol? sqrt(0) = 0

OpenStudy (anonymous):

well if you feel more confortable with exponentials: sqrt(0) = 0^(1/2) = 0

OpenStudy (anonymous):

\[\sqrt{0} = \sqrt{0^2} = 0\]